OpenStudy (anonymous):

The sales tax in one state is 7%. Write a function rule for finding the total cost of an item with selling price x. Then find the total cost of a CD player with a selling price of $150.

OpenStudy (anonymous):

a. f(x) = x – 7; $148.93 c. f(x) = 7 + x; $151.07 b. f(x) = 0.07x; $10.50 d. f(x) =1.07x; $160.50

OpenStudy (anonymous):

I need help!!!!!